0.25 mol of an ideal monoatomic gas undergoes isothermal expansion from a
volume of 2.0 dm3
to 10 dm3
at 27 oC. Calculate the maximum work that can be
obtained from this process.
1
Expert's answer
2020-02-27T09:04:27-0500

Wmax = nRTln(V2/V1) = 0.25 * 8.314 * (27+273) * ln(10/2) = 1004 J